OpenStudy (anonymous):

Given: Square with side c. All four interior triangles are right triangles. All four interior triangles are congruent. The interior quadrilateral is a square. Prove: a2 + b2 = c2

OpenStudy (anonymous):

When written in the correct order, the paragraph below proves the Pythagorean Theorem using the diagram. Let a represent the height and b represent the base of each triangle. The area of one triangle is represented by the expression One halfab. (1) The length of a side of the interior square is (a – b). (2) The area of all four triangles will be represented by 4 • One halfab or 2ab. (3) The area of the interior square is (a – b)2. (4) By distribution, the area is a2 – 2ab + b2. The area of the exterior square is found by squaring side c, which is c2, or by adding the areas of the four interior triangles and interior square, 2ab + a2 – 2ab + b2. Therefore, c2 = 2ab + a2 – 2ab + b2. Through addition, c2 = a2 + b2.

OpenStudy (anonymous):

Which is the most logical order of statements (1), (2), (3), (4) to complete the proof? (2), (1), (4), (3) (1), (2), (3), (4) (1), (2), (4), (3) (2), (1), (3), (4)

OpenStudy (anonymous):

hey king geogre

OpenStudy (anonymous):

this is another geometry problem that has been giving me trouble :(

OpenStudy (anonymous):

you think that you can help?

OpenStudy (anonymous):

hey guys

OpenStudy (anonymous):

this is another tough geometry problem

OpenStudy (anonymous):

you think you guys can help me with this challenge?

OpenStudy (kinggeorge):

First off, which order do you think is correct and why?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

One seems to go first

OpenStudy (anonymous):

so that eliminates choices 1 and four

OpenStudy (kinggeorge):

I agree so far.

OpenStudy (anonymous):

so 2 would have to go next

OpenStudy (anonymous):

then 3 would have to go third

OpenStudy (anonymous):

and 4 sums up the areas

OpenStudy (kinggeorge):

Perfect.

OpenStudy (anonymous):

wow...

OpenStudy (anonymous):

That wasn't so hard when you guided me...

OpenStudy (kinggeorge):

Exactly. Most of these you just need a push in the right direction.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

and the duplicate I chose for question 4 on that last assignment was incorrect

OpenStudy (anonymous):

remember the question with two identical answer choices?

OpenStudy (kinggeorge):

yup.

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

hmm...

OpenStudy (kinggeorge):

Which did you guess?

OpenStudy (anonymous):

I guessed the third one on the question 4 thing

OpenStudy (anonymous):

I guess I answered the wrong duplicate :(

OpenStudy (kinggeorge):

sad day :(

OpenStudy (anonymous):

I need to ask you something

OpenStudy (anonymous):

is f squared= a squared + b squared an example of the reflexive property of equality?

OpenStudy (kinggeorge):

No. I wouldn't think so.

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

a=a

OpenStudy (anonymous):

f squared= a squared +b squared

OpenStudy (kinggeorge):

a=a would be an example.

OpenStudy (anonymous):

3=3

OpenStudy (anonymous):

25=9+16

OpenStudy (anonymous):

25=25?

OpenStudy (anonymous):

It would not be substitution, right?

OpenStudy (kinggeorge):

anything that's an example of reflexive is something that looks like \(a=a\). Which of these look like that?

OpenStudy (anonymous):

well...

OpenStudy (anonymous):

not f squared=a squared plus b squared

OpenStudy (anonymous):

could it be substitution instead?

OpenStudy (kinggeorge):

I'm not sure what that one is without more context. However, substitution sounds like it might be correct.

OpenStudy (anonymous):

whats an example of substitution?

OpenStudy (anonymous):

(A) + B = 12 (2B) + B = 12 3B = 12 B = 4

OpenStudy (anonymous):

thats one example

OpenStudy (kinggeorge):

\(a=g+c\), \(g=h \implies a=h+c\). Is one example. What you gave is another example.

OpenStudy (anonymous):

so f squared= a squared plus b squared must be an eaxample, right?

OpenStudy (kinggeorge):

What are the other options?

OpenStudy (anonymous):

Reason #1 - Reflexive Property of Equality Reason #2 - SSS Postulate Reason #1 - Substitution Reason #2 - SAS Postulate Reason #1 - Substitution Reason #2 - SSS Postulate Reason #1 - Reflexive Property of Equality Reason #2 - SAS Postulate

OpenStudy (anonymous):

I'm thinking the second choice

OpenStudy (anonymous):

they are trying to prove that a triangle has a right angle

OpenStudy (anonymous):

thats why I go for SAS

OpenStudy (anonymous):

Given: In ∆ACB, c2 = a2 + b2. Prove: ∆ACB is a right angle. Complete the flow chart proof with missing reasons to prove that ∆ACB is a right angle.

OpenStudy (anonymous):

Triangles ACB and DFE are shown below.

OpenStudy (anonymous):

anything?

OpenStudy (kinggeorge):

First one, reason 1 is definitely substitution. However, I would use SSS instead SAS because you're given 3 sides, and not angles.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So you are thinking choice three?

OpenStudy (kinggeorge):

That's what I would choose.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ty

OpenStudy (kinggeorge):

you're welcome.

OpenStudy (anonymous):

ehhh

OpenStudy (anonymous):

can you help me with the last geometry problem of the day?

OpenStudy (anonymous):

just one more...

OpenStudy (anonymous):

just one more...

OpenStudy (anonymous):

Given: In ∆ABC, segment DE is parallel to segment AC . Prove: BD over BA equals BE over BC

OpenStudy (anonymous):

The two-column proof with missing statements and reasons proves that if a line parallel to one side of a triangle also intersects the other two sides, the line divides the sides proportionally.

OpenStudy (anonymous):

Statement 1: Line segment DE is parallel to line segment AC

OpenStudy (anonymous):

Reason is given

OpenStudy (anonymous):

Statement 2: Line segment BA is a transversal that intersects two parallel lines

OpenStudy (anonymous):

Reason: Conclusion from statement 1

OpenStudy (anonymous):

3. Statement and Reason blank

OpenStudy (anonymous):

Angle B is congruent to Angle B(4.)

OpenStudy (anonymous):

Reason: Reflexive property of equality

OpenStudy (anonymous):

Triangle ABC is similar to triangle DBE

OpenStudy (anonymous):

Reason: Angle-Angle(AA) Similarity Postulate

OpenStudy (anonymous):

Statement and Reason for 6. is unknown

OpenStudy (anonymous):

Complete the proof by entering the correct statements and reasons.

OpenStudy (anonymous):

hey agent

OpenStudy (kinggeorge):

Are there any options for the statements/reasons?

OpenStudy (anonymous):

sadly no :(

OpenStudy (anonymous):

I have to fill them in

OpenStudy (anonymous):

any suggestions or tips?

OpenStudy (anonymous):

what say you monkey?

OpenStudy (kinggeorge):

for 6, it's pretty easy. It's the very last statement in the proof. What does this mean it has to be?

OpenStudy (anonymous):

BD over BA equals BE over BC

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

and the reason is

OpenStudy (anonymous):

hmm...